OpenStudy (anonymous):

find the limit [(sin^2)x]/tanx as x approaches 0.

OpenStudy (anonymous):

cos^2x=1

OpenStudy (anonymous):

whoops didnt see the tan x

OpenStudy (anonymous):

rewrite as sin^2(x)/[sinx/cosx] = sin(x)cos(x) The limit of sin(x)cos(x) can be thought of as the product of the limit of sin(x) and cos(x) separately. sin(x) approaches 0 as x apparoaches 0, and cos(x) approaches 1 as x apparoches 0, so the limit of sin(x)cos(x) = 0(1) = 0

OpenStudy (anonymous):

sin^2x/tanx = cos2x/sec^2x= cos^2x*cos^2x= 1 using L' Hospital's rule. Can you use that rule pyeh?

OpenStudy (anonymous):

Yes you can, but it's easier to rewrite.

OpenStudy (anonymous):

Also, derivative of sin^2(x) = 2sin(x)cos(x)

OpenStudy (anonymous):

why did i get 1 using that rule, and you got 0?

OpenStudy (anonymous):

see the above post; I think you took the derivative of the numerator incorrectly.